OpenStudy (anonymous):

(4/31/20/3)-4/25*20/3=

OpenStudy (shadowfiend):

So you would just do this left to right. 4/31 / 20 is the same as 4/ (31 * 20), which is 4 / 620. 4/620 / 3 is the same as 4 / (620 * 3) = 4/ 1860. So then you have 4 / 1860 - 4 / 25 * 20 / 3. Order of operations tells us to do the multiplication/division first, so we can multiply 4/25 and 20/3 together to get 80/75. So you have 4 / 1860 - 80/75. You now need to find a common denominator. First, though, we'll simplify the first fraction by dividing the top and bottom by 4 and get 1 / 465 - 80/75. We can then multiply 75 * 465 to get our common denominator of 34875. We multiply 1 / 465 by 75/75 to get 75/34875 and we multiply 80/75 by 465/465 to get 37200/34875. Finally, we have 75/34875 - 37200/34875 = -37125/34875. It turns out, both of these can be divided evenly by 225, so we can simplify to 165 / 155, and then simplify again by dividing by 5 to get 33/31.